OpenStudy (anonymous):

For a certain bathtub, it takes the hot water faucet 3 times as long to fill the tub as it does the cold water faucet. Left on together, the two faucets take 6 minutes to fill the tub. How long will it take the hot water faucet to fill the tub by itself?

OpenStudy (gorv):

let cold faucet take x minute hot water faucet will take =3x minute 1/x+1/3x=1/6 (3x+x)/3x^2=1/6 4x/3x^2=1/6 4/3x=1/6 4*6=3x 24=3x x=24/3=8

OpenStudy (anonymous):

It says its wrong

OpenStudy (gorv):

loll x= cold water faucet hot water faucet=3*x=3*8=24